chicken egg mas the mass of 60kg.The egg is 10% of the total mass.The egg shall is made of calcium carbonate what is the mass of the calcium in the egg


1
Expert's answer
2022-03-27T10:27:14-0400

The egg shell has mass 100-10=90%

0.9x60kg =5.4 kg of calcium carbonate.

M(CaCO3)=40+12+16x3=100 g/mol

n(CaCO3)= 5,4x1000/100=54 mole

M(Ca)=40 g/ mol

n(Ca)=n(CaCO3)

m(Ca)=40x54=2160 g